An object of mass 0.50 kg weighs 2.0 N when it is 1.0 × 106 m above the surface of Planet X. The radius of the planet is 4.0 × 106 m. What free-fall acceleration will the 0.50-kg object experience when at the surface of Planet X?

A) 2.0 m/s2

B) 6.3 m/s2

C) 4.0 m/s2

D) 4.8 m/s2
Answer #1

Gravitational potential, g' = g/(1 + (h/R))^2

g' = 2/0.5 = 4 m/s^2

h = 1 x 10^6 m

R = 4 x 10^6 m

Substituting the values,

4 = g/(1 + (1/4))^2

g = 6.3 m/s^2

Comment in case any doubt please rate my answer....